OpenStudy (anonymous):

Evaluate the following expression for x = –2 and y = –4. 4(x+y)^4 A. –64 B. 64 C. –5,184 D. 5,184

OpenStudy (anonymous):

i think is b but im not sure

OpenStudy (mchilds15):

-2+-4=?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

FIrst add -2 + -4

OpenStudy (mchilds15):

then multiply times 4 ... and then you do ^4

OpenStudy (anonymous):

Can you add -2 + -4? @rocky1718

OpenStudy (anonymous):

so it will be -5,184 ?

OpenStudy (anonymous):

Yes, but there is no negative.

OpenStudy (anonymous):

why

OpenStudy (anonymous):

-6 * -6 * -6 * -6 = 1296 1296 * 4 = 5,184

OpenStudy (mchilds15):

^

OpenStudy (anonymous):

A negative times a negative equals a positive. -6 * -6 = 36 36 * -6 = -216 -216 * -6 = 1296

OpenStudy (anonymous):

mmok thx @iGreen.